# FaceLift: Single Image to 3D Head with View Generation and GS-LRM

### v1

@@ Abstract
We present FaceLift, a feed-forward approach for rapid, high-quality 360-degree head reconstruction from a single image. Our pipeline begins by employing a multi-view latent diffusion model that generates consistent side and back views of the head from a single facial input. These generated views then serve as input to a GS-LRM reconstructor, which produces a comprehensive 3D representation using Gaussian splats. To train our system, we develop a dataset of multi-view renderings using synthetic 3D human head assets. The diffusion-based multi-view generator is trained exclusively on synthetic head images, while the GS-LRM reconstructor undergoes initial training on Objaverse followed by fine-tuning on synthetic head data. FaceLift excels at preserving identity and maintaining view consistency across views. Despite being trained solely on synthetic data, FaceLift demonstrates remarkable generalization to real-world images. Through extensive qualitative and quantitative evaluations, we show that FaceLift outperforms state-of-the-art methods in 3D head reconstruction, highlighting its practical applicability and robust performance on real-world images. In addition to single image reconstruction, FaceLift supports video inputs for 4D novel view synthesis and seamlessly integrates with 2D reanimation techniques to enable 3D facial animation. Project page: https://weijielyu.github.io/FaceLift.
Figure 1: FaceLift takes a single image of a human face as input and generates a high-fidelity 3D Gaussian head representation. We present the rendering result on the left side and the centers of the Gaussians on the right side for visualization. The generated Gaussian representation enables high-quality, full-head novel view synthesis (NVS) while accurately capturing fine details of the face and hair.
@@ 1 Introduction
3D head reconstruction has been a central focus in computer vision and computer graphics research for decades, driven by its crucial applications in virtual and augmented reality, digital entertainment, and telepresence systems. The task remains particularly challenging because the human visual system is highly attuned to facial details, making even subtle rendering artifacts far more noticeable than those in other objects.
@@
Traditional approaches to 3D head synthesis typically use parametric textured mesh models [54, 27] trained on large 3D scan datasets. While these models enable basic head generation, the resulting rendered images often lack fine-scale geometric and textural details, which limits their perceptual realism and expressive capabilities. Recent breakthroughs in image generative models [15, 19] and novel view synthesis techniques [35, 23] have opened new possibilities for this research area. Leveraging these developments, recent works [62, 1] use neural 3D representations and large, unstructured face image datasets to learn effective 3D head representations without requiring explicit multi-view consistency in the training data. A parallel approach such as RodinHD [63] leverages synthetic multi-view images to train generative models that directly output 3D neural representations of the head. However, training solely on synthetic data often results in significant perceptual identity loss in the generated outputs, as demonstrated in Fig. 2.
@@
Figure 2: RodinHD [63] trains triplane diffusion with only synthetic data, resulting in apparent identity loss. By leveraging stable diffusion priors [42], FaceLift achieves better identity preservation and generalizes effectively to real human portraits.
Diffusion-based view generation methods have achieved remarkable results in text-to-3D and image-to-3D tasks [31, 47, 32, 26, 25]. These approaches leverage text-to-image foundation models trained on vast internet data, combined with training on synthetic 3D object renderings [8, 9], demonstrating impressive generalization capabilities for general object view synthesis. However, due to the high perceptual sensitivity of human faces, we show that models trained solely on object data fail to generalize effectively to 3D human face reconstruction.
Inspired by this, we present FaceLift, a two-stage pipeline for high-quality 3D head reconstruction from a single image. The first stage builds upon diffusion-based multi-view generation methods [32, 26], leveraging an image-conditioned diffusion model [42] as the core view generation backbone. By fine-tuning this model on synthetic human head data, we achieve two key advantages: (1) robust view consistency, enabled by strong priors from multi-view consistent training data, and (2) enhanced generalization capability, inherited from the large foundational model, ensuring accurate identity preservation during inference. In the second stage, we employ a state-of-the-art large reconstruction model [64] to fuse the generated sparse views into a comprehensive 3D Gaussian representation. Specifically, we fine-tune GS-LRM using synthetic human head data, enabling the reconstruction of detailed facial geometry and appearance.
@@
We evaluate FaceLift both quantitatively and qualitatively across diverse datasets. Using real multi-view studio captures [34] and an independent synthetic human dataset [6], our approach consistently surpasses previous state-of-the-art methods across all evaluation metrics. Through extensive testing on in-the-wild portrait images, we demonstrate that FaceLift reconstructs complete 3D heads with (1) fine-grained geometric and appearance details, (2) accurate identity preservation, and (3) high visual fidelity. Through comparisons with general object 3D reconstruction models and detailed ablation studies, we demonstrate that multi-view consistent training data plays a crucial role in achieving high-fidelity face reconstruction results.
Our contributions are summarized as follows:
@@
  • We propose FaceLift, a two-stage framework that reconstructs a high-fidelity 3D head from a single image using view generation and large reconstruction model.
  • Our approach combines synthetic human head data for view-consistent training with the rich priors from a large image generation foundation model, enabling robust generalization capabilities.
@@
  • Through comprehensive quantitative and qualitative evaluation, we demonstrate that our approach achieves state-of-the-art performance in reconstruction accuracy and identity preservation.
@@ 2 Related Work
Face Reconstruction. 3D face reconstruction has long been a long-standing challenge in computer vision, with substantial progress driven by diverse approaches. Vetter and Blanz [54] pioneer a method for synthesizing 3D faces by linearly blending multiple 3D templates, now widely known as blendshapes. This work establishes the foundation for 3D Morphable Models (3DMMs), which represent 3D face shapes and textures as principal components derived from scanned data. Subsequent research [3, 4, 27, 29, 41] extend this framework, enabling the generation of new 3D faces by manipulating blending coefficients. However, these methods produce mesh-based representations that lack fine details and are limited to modeling the front of the face, excluding hair and 360-degree synthesis.
@@
While 3DMM-based methods have been foundational, recent advances in deep learning, especially Generative Adversarial Networks (GANs) [15, 21, 22], have greatly improved 3D face synthesis quality. EG3D [62] uses a tri-plane NeRF representation with a pose-conditioned StyleGAN2 [22] framework. Follow-up works [28, 2] achieve single-image-to-3D generation through GAN inversion [7]. Despite their success, these methods can only synthesize near-frontal views. To overcome this, PanoHead [1] introduces a tri-grid neural volume representation, enabling full 360-degree head synthesis. Unfortunately, it does not provide a 3D head representation for consistent multi-view rendering. Recent efforts explore alternative representations for 3D face reconstruction from sparse input, such as a single image [13, 37, 55] or few-shot images [5]. However, these methods still require pre-instance optimization. Rodin [57] and its extension RodinHD [63] employ an image-conditioned diffusion model to generate a triplane representation of a human head for full-head novel view synthesis. Nevertheless, their triplane diffusion model is limited to synthetic data and struggles to achieve high-fidelity reconstructions from real-world images.
@@
Synthetic Human Data. Capturing high-quality 3D data of real humans requires a controlled studio environment and costly photography equipment [34]. As an alternative, large-scale synthetic 3D head datasets have emerged as an effective and resource-efficient solution for tasks like human head reconstruction [59, 57, 63, 6] and photorealistic relighting [61], offering a scalable way to train models without the restrictions of real-world data acquisition. Inspired by these previous works, we aim to use synthetic data to teach the model knowledge of the human head and minimize the generalization gap between synthetic data training and real-world inference.
@@
Image or Text to 3D. Generative models have achieved remarkable success in 2D image generation with VAEs [24, 52], GANs [15, 21, 22], and diffusion models [19, 42, 48]. Building on this success, extensive research has extended these models to 3D content generation [60, 14, 38, 36]. Starting with DreamFusion [39], numerous works [31, 40, 50, 58, 45] try to distill NeRF [35, 62] or 3D Gaussians [23] representation from 2D image diffusion using a Score Distillation Sampling (SDS) loss. These methods can produce high-quality results but often encounter challenges such as slow optimization, over-saturated colors, and the Janus problem. To overcome these challenges, recent works [30, 47, 32, 26, 25] generate multi-view images with high consistency, which can be directly used for 3D reconstruction with neural reconstruction methods [35, 23, 56]. However, optimizing NeRF or NeuS remains far from real-time performance.
Recent advancements in large reconstruction models (LRMs) [20, 25, 64, 51] offer a pathway to faster 3D reconstruction. Leveraging scalable transformer architectures [53, 11] and large datasets like Objaverse [8, 9], these models effectively capture generalizable 3D priors. Unlike traditional pre-scene optimization methods [35, 23, 56], LRMs employ a feed-forward approach, enabling the prediction of high-quality NeRF, mesh, or 3D Gaussian representations from sparse images in under a second. However, most of these research efforts are applied to general objects, with limited or suboptimal results demonstrating their ability on 3D head reconstruction.
@@
Figure 3: Overview of FaceLift. Given a single image of a human face as input, we train an image-conditioned, multi-view diffusion model to generate novel views covering the entire head. By leveraging pre-trained weights and high-quality synthetic data, our multi-view latent diffusion model can hallucinate unseen views of the human head with high-fidelity and multi-view consistency. We then fine-tune a GS-LRM [64], which takes multi-view images and their camera poses as input and generates 3D Gaussian splats to represent the human head. The generated 3D Gaussian representation enables full-head novel view synthesis.
@@ 3 Proposed Method
Given a single frontal image of a human face, our goal is to reconstruct a complete 3D head in the form of 3D Gaussian splats. We aim to generate plausible details not visible in the input view, such as the back of the head, while ensuring coherent visualization from any angle. Our proposed FaceLift follows a two-stage pipeline, as shown in Fig 3. In the first stage, using multi-view synthetic images of human heads, we train an image-conditioned diffusion model to generate six views of a human head, maintaining multi-view consistency and preserving identity. These generated views, in the second stage, along with their corresponding camera poses, are then input into a fine-tuned GS-LRM model to reconstruct a set of 3D Gaussian splats.
@@ 3.1 Synthetic Human Head Dataset
Figure 4: Synthetic data examples. Top: six views for view generation. Bottom: samples of random views for GS-LRM training.
@@
To train our networks, we render images (samples shown in Fig. 4) at 512$\times$512 resolution from 200 unique identities, each with 50 varied appearances, including different hairstyles, skin tones, expressions, clothes, poses, etc. We render our training dataset under two types of lighting conditions: (1) ambient light and (2) random HDR environment light. We render six views for each subject to train the multi-view diffusion model. For fine-tuning GS-LRM [64], we render 32 views with random camera poses.
3.2 Stage I: Single Image to Multi-view Generation
Recent 3D generation and reconstruction frameworks [47, 32, 26, 25] have demonstrated promising results by leveraging diffusion models to generate multi-view consistent images. Building on this approach, we adopt the open-source Stable Diffusion text-to-image generator [42] as the foundation of our view generation network. Specifically, we utilize the Stable Diffusion V2-1-unCLIP model [43], which is fine-tuned on Stable Diffusion V2-1 to incorporate CLIP image embeddings as additional conditioning input. In our pipeline, a CLIP image encoder processes the input frontal view face images, generating embeddings that serve as conditioning signals for the diffusion model.
To ensure the consistency of the generated novel views, we adopt a multi-view attention mechanism similar to [32, 26] to facilitate information propagation and implicitly encode multi-view dependencies. The diffusion model learns multi-view correlations by sharing information across views within the attention layers, allowing it to generate consistent RGB images from multiple perspectives.
Given a single near frontal view face image with azimuth $\alpha$, the multi-view diffusion model will generate six views with azimuths equal to $\alpha$, $\alpha\pm 45^{\circ}$, $\alpha\pm 90^{\circ}$, $\alpha+180^{\circ}$}, covering 360 degrees of the human head. All images, both input and generated output, maintain a zero elevation angle, ensuring consistent horizontal viewpoints. The generated views consist of a reconstructed front view matching the input image, left and right profiles capturing the sides of the head, and a back view that synthesizes hair structure and color based on the frontal input and learned priors. We also generate three-quarter views (left-front and right-front) to enhance facial details in the following reconstruction stage.
3.3 Stage II: Multi-view to 3D Gaussian Head Reconstruction
Stage II conducts sparse view reconstruction using the six views generated by Stage I. For this task, we utilize GS-LRM [64], a state-of-the-art reconstruction model that excels at generating detailed 3D models from limited viewpoints. GS-LRM’s use of 3D Gaussians as its underlying representation is particularly effective for capturing the intricate details of human heads, such as hair strands. Furthermore, its feed-forward architecture enables rapid reconstruction, making it well-suited for practical applications.
GS-LRM employs a transformer architecture to regress pixel-aligned 3D Gaussians from a set of posed images. Each of the multi-view images is concatenated with its corresponding Plücker ray coordinates derived from the camera’s intrinsic and extrinsic parameters for pose conditioning. The images are divided into non-overlapping patches, with each 2D patch flattened into a 1D vector. A linear layer then maps these vectors to image patch tokens. The set of multi-view image tokens is concatenated and processed through a chain of transformer blocks. Later, the output tokens from the transformer are decoded into Gaussian parameters using a single linear layer. Then, the Gaussian parameters are unpatchified into Gaussians. Each 2D pixel corresponds to one 3D Gaussian. Gaussian is parameterized by 3-channel RGB, 3-channel scale, 4-channel rotation quaternion, 1-channel opacity, and 1-channel ray distance, similar to [23]. The final output of the GS-LRM model is the merge of 3D Gaussians from all input views.
During GS-LRM training, we randomly select four views from the available camera poses to reconstruct a total of eight views—four input views plus four novel viewpoints. Following [64], we optimize the model using a combination of MSE and perceptual losses. At inference time, GS-LRM processes the outputs from our multi-view diffusion model. While GS-LRM can accommodate any number of input views, we utilize all six views generated by the diffusion model to achieve optimal visual quality.

### v2

# FaceLift: Learning Generalizable Single Image 3D Face Reconstruction from Synthetic Heads

@@ Abstract
We present FaceLift, a novel feed-forward approach for generalizable high-quality 360-degree 3D head reconstruction from a single image. Our pipeline first employs a multi-view latent diffusion model to generate consistent side and back views from a single facial input, which then feed into a transformer-based reconstructor that produces a comprehensive 3D Gaussian splats representation. Previous methods for monocular 3D face reconstruction often lack full view coverage or view consistency due to insufficient multi-view supervision. We address this by creating a high-quality synthetic head dataset that enables consistent supervision across viewpoints. To bridge the domain gap between synthetic training data and real-world images, we propose a simple yet effective technique that ensures the view generation process maintains fidelity to the input by learning to reconstruct the input image alongside the view generation. Despite being trained exclusively on synthetic data, our method demonstrates remarkable generalization to real-world images. Through extensive qualitative and quantitative evaluations, we show that FaceLift outperforms state-of-the-art 3D face reconstruction methods on identity preservation, detail recovery and rendering quality.
Figure 1: FaceLift transforms a single facial image into a high-fidelity 3D Gaussian head representation. Trained exclusively on synthetic 3D data, our pipeline first generates sparse, identity-preserving multiview images of the input head using a diffusion model. These sparse generated views are then fed into a transformer-based 3D Gaussian splats reconstructor, producing complete and detailed 3D head representation that generalize remarkably well to real-world human images. Project page: https://weijielyu.github.io/FaceLift.
@@ 1 Introduction
3D face reconstruction has been a central focus in computer vision and graphics research for decades, driven by its crucial applications in immersive virtual and augmented realities, VFX and gaming, digital entertainment, and next-generation telepresence systems. However, achieving high quality reconstruction from a single image remains very challenging. On one hand, the monocular face reconstruction problem is highly ill-posed – a single 2D image can be produced by countless different 3D face shapes, creating fundamental ambiguity. On the other hand, the human visual system is highly attuned to facial details, making even subtle artifacts and imperfections noticeable to the eye.
@@
Traditional 3D head synthesis approaches typically use parametric textured mesh models [60, 32] trained on 3D scan datasets. While these models enable basic head generation, the rendered images frequently lack fine-scale geometric detail, realistic textures, and convincing hair, limiting their perceptual realism and expressiveness. Recent breakthroughs in image generative models [19, 23] and novel view synthesis techniques [40, 27] have opened new possibilities for this research area. Leveraging these developments, recent works [72, 1] use neural 3D representations to learn effective 3D head representation from unstructured real face image datasets [25, 68]. While these datasets improve the realism and diversity of rendering results, they fail to provide multi-view supervision for modeling 3D consistency causing view inconsistency and artifacts on the back of the head. Since diverse multi-view real images are hard to acquire, RodinHD [73] leverages synthetic multi-view images to train generative models that directly output 3D neural representations of the head. However, training solely on synthetic data often results in significant perceptual identity loss in the generated outputs, as demonstrated in Fig. 2.
@@
Figure 2: Comparison with RodinHD. RodinHD [73] trains triplane diffusion with synthetic data, resulting in apparent identity loss. In contrast, FaceLift achieves better identity preservation and generalizes effectively to real human portraits.
In this work, we present FaceLift, a pipeline for learning generalizable and high-fidelity single image to 3D face representation from synthetic head data. To achieve high quality reconstruction that preserves the input facial identity, we adopt a two-stage pipeline to first generate identity preserving multi-view images using a diffusion model [48], followed by a transformer-based feed-forward reconstructor to fuse the generated sparse views into a comprehensive 3D Gaussian representation. We train the model with synthetic images – multi-view renderings of 3D synthetic human portraits using Blender. We highlight two key techniques for generalizing to real-world images and preserving input facial identity. First, we emphasize the importance of reconstructing the input image alongside the view synthesis task in the conditional diffusion model training, which significantly improved generalization capability in testing. Second, we demonstrate that training the feed-forward reconstructor benefits from a two-stage training process: pre-training on general objects [12] to acquire a rich geometry and texture prior, followed by fine-tuning on synthetic human head data to capture head-specific geometry. With our two-stage approach, we focus on learning identity preservation in the image space during the first stage, achieving higher input fidelity compared to existing methods.
Comparing with prior art, we achieve three key advancements: (1) robust view consistency through multi-view attention and supervision, (2) improved generalization from training techniques and foundational model, ensuring accurate identity preservation, (3) high-quality facial texture and hair details via pixel-aligned Gaussian representation.
@@
We extensively evaluate FaceLift quantitatively and qualitatively across diverse datasets. Using real multi-view studio captures [39] and an independent synthetic human dataset [8], our approach consistently surpasses previous state-of-the-art methods across all evaluation metrics. Through extensive testing on in-the-wild portrait images, we demonstrate that our method reconstructs complete 3D heads with fine-grained details, accurate identity preservation, and high visual fidelity. Comparisons and ablation studies confirm that multi-view consistent training data is crucial for high-fidelity face reconstruction. Our contributions are summarized as follows:
@@
  • We propose FaceLift, a framework that reconstructs a high-fidelity 3D head from a single image using view generation and feed-forward reconstructor.
  • Despite being trained solely on synthetic human head data, our approach shows no domain gap on real-world images, highlighting both the effectiveness of synthetic data and our model’s robust generalization capabilities.
  • We construct two benchmarks on single-image to 3D full head reconstruction tasks using the publicly available datasets Cafca [8] and Ava-256 [39] to quantitatively evaluate models’ performance on both reconstruction accuracy and identity preservation ability.
@@
  • Our comprehensive evaluation confirms that our approach achieves state-of-the-art performance in reconstruction accuracy and identity preservation.
@@ 2 Related Work
Face Reconstruction. 3D face reconstruction has been a long-standing challenge in computer vision, with substantial progress driven by diverse approaches. Vetter and Blanz [60] pioneer a method for synthesizing 3D faces by linearly blending multiple 3D templates, now widely known as blendshapes. This work establishes the foundation for 3D Morphable Models (3DMMs), which represent 3D face shapes and textures as principal components derived from scanned data. Subsequent research [5, 6, 32, 34, 47] extend this framework, enabling the generation of new 3D faces by manipulating blending coefficients. However, these methods produce mesh-based representations that lack fine details and are limited to modeling the front of the face, excluding hair and 360-degree synthesis. While 3DMM-based methods have been foundational, recent advances in deep learning, especially Generative Adversarial Networks (GANs) [19, 25, 26], have greatly improved 3D face synthesis quality. EG3D [72] uses a tri-plane NeRF representation with a pose-conditioned StyleGAN2 [26] framework. Follow-up works [33, 3] achieve single-image-to-3D generation through GAN inversion [11]. Despite their success, these methods can only synthesize near-frontal views. To overcome this, PanoHead [1] introduces a tri-grid neural volume representation, enabling full 360-degree head synthesis. Unfortunately, it does not provide a 3D head representation for consistent multi-view rendering. Recent efforts explore alternative representations for 3D face reconstruction from sparse input, such as a single image [17, 42, 61] or few-shot images [7]. However, these methods still require pre-instance optimization. Rodin [63] and its extension RodinHD [73] employ an image-conditioned diffusion model to generate a triplane representation of a human head for full-head novel view synthesis. Nevertheless, their triplane diffusion model is limited to synthetic data and struggles to achieve high-fidelity reconstructions from real-world images. For animatable 3D head avatars generations, Morphable Diffusion [10] generates multi-view consistent images from a single image using a morphable mesh, while HeadGAP [76] generates 3D animatable head avatars using few-shot input, leveraging 3D head priors learned from large-scale data. In contrast, our work focuses on leveraging synthetic training data to produce high-fidelity, detailed 3D Gaussian head models.
@@
Synthetic Human Data. Capturing high-quality 3D data of real humans requires a controlled studio environment and costly photography equipment [39]. As an alternative, large-scale synthetic 3D head datasets have emerged as an effective and resource-efficient solution for tasks like human head reconstruction [65, 63, 73, 8] and photorealistic relighting [71, 9], offering a scalable way to train models without the restrictions of real-world data acquisition. Inspired by these previous works, we aim to use synthetic data to improve the model’s understanding of the human head and minimize the generalization gap between synthetic data training and real-world inference.
@@
Image or Text to 3D. Generative models have achieved remarkable success in 2D image generation with VAEs [28, 58], GANs [19, 25, 26], and diffusion models [23, 48, 54]. Building on this success, extensive research has extended these models to 3D content generation [66, 18, 43, 41]. Starting with DreamFusion [45], numerous works [36, 46, 57, 64, 51] try to distill NeRF [40, 72, 62] or 3D Gaussians [27] representation from 2D image diffusion using a Score Distillation Sampling (SDS) loss. These methods can produce high-quality results but often encounter challenges such as slow optimization, over-saturated colors, and the Janus problem. To overcome these challenges, recent works [35, 53, 37, 31, 30] generate multi-view images with high consistency, which can be directly used for 3D reconstruction with neural reconstruction methods [40, 27, 62]. However, optimizing NeRF or NeuS remains far from real-time performance. Recent advancements in large 3D reconstruction models (LRMs) [24, 30, 74, 56] offer a pathway to faster 3D reconstruction. Leveraging scalable transformer architectures [59, 15] and large datasets like Objaverse [12, 13], these models effectively capture generalizable 3D priors. Unlike traditional pre-scene optimization methods [40, 27, 62], LRMs employ a feed-forward approach, enabling the prediction of high-quality NeRF, mesh, or 3D Gaussian representations from sparse images in under a second. However, most of these research efforts are applied to general objects, with limited or suboptimal results for 3D head reconstruction.
@@
Figure 3: Overview of FaceLift. Given a single image of a human face $y$ as input, we train an image-conditioned, multi-view diffusion model to generate novel views $x_{0}^{1},\dots,x_{0}^{N}$ covering the entire head. By generating $x_{0}^{1}$ the same as $y$ and leveraging the high-quality synthetic data, our multi-view latent diffusion model can hallucinate unseen views of the human head with high-fidelity and multi-view consistency. We then train a transformer-based reconstructor $f_{R}$, which takes multi-view images $x_{0}^{1:N}$ and their camera poses $\textsc{P}^{1:N}$ as input and generates 3D Gaussian splats $G$ to represent the human head.
@@ 3 Proposed Method
As shown in Fig. 3, given a single frontal image of a human face $y$, our goal is to reconstruct a complete 3D head $G$, represented as 3D Gaussian splats, with detailed facial texture and preserved identity. This requires our system to have prior knowledge on the geometry structure of a human face and the ability to synthesis plausible details which are not visible in the input view. Hence, we train a multi-view diffusion model $f_{D}$ on synthetic human head data to generate $N$ views $x_{0}^{1},x_{0}^{2},\dots,x_{0}^{N}$ covering $360^{\circ}$ of the human head while achieving multi-view consistency and preserving identity. We choose pixel-aligned 3D Gaussians to obtain the final 3D representation. Compared to NeRFs and meshes, 3D Gaussians offer explicit volumetric primitives that better capture subtle facial geometry and fine details. Their semi-transparent kernels naturally model effects like wispy hair and translucency, which are challenging for discrete surfaces or density fields. These generated views $x_{0}^{1:N}$ from the diffusion model, along with their corresponding Plücker ray coordinates $\textsc{P}^{(1:N)}$, are input into a transformer-based Gaussian reconstructor $f_{R}$ to predict a set of 3D $G$. Training of the Gaussian reconstructor follows a pre-training process on general objects [12] and a fine-tuning process on synthetic human head data.
@@ 3.1 Synthetic Human Head Dataset
Figure 4: Synthetic data examples. Top row: six views for diffusion training. Bottom row: samples of random views for reconstructor training.
@@
To train our networks, we render images (samples shown in Fig. 4) at 512$\times$512 resolution from 200 unique identities, each with 50 varied appearances, including different hairstyles, skin tones, expressions, clothes, poses, etc. We render our training dataset under two types of lighting conditions: (1) ambient light and (2) random HDR environment light. We render six views for each subject to train the multi-view diffusion model. For fine-tuning the transformer-based reconstructor, we render 32 views with random camera poses.
3.2 View Generation
We model the sparse view generation from a single image input as a conditional diffusion process. We use a multi-view diffusion model $f_{D}$ to generate $N$ views, denoted as $X_{0}^{1},X_{0}^{2},\dots,X_{0}^{N}$, given a single front-facing image $y$ and CLIP text embeddings $e^{1},e^{2},\dots,e^{N}$ corresponding to each generated view. This process is expressed as:
$\{X_{0}^{1},X_{0}^{2},\dots,X_{0}^{N}\}=f_{D}(y,\{e^{1},e^{2},\dots,e^{N}\}).$ (1)
We aim to learn the joint distribution of all these views, conditioning on the input image $y$ and text embedding $e^{1},e^{2},\dots,e^{N}$. We denote the joint distribution as:
$p_{f_{D}}(x_{0}^{1:N}\mid y,e^{1:N}):=p_{f_{D}}(\{x_{0}^{1},\dots x_{0}^{N}\}\mid y,\{e^{1},\dots e^{N}\}).$ (2)
View Selection. Given a single near frontal view face image with azimuth $\alpha$, our multi-view diffusion model generates six views with azimuths equal to $\alpha$, $\alpha\pm 45^{\circ}$, $\alpha\pm 90^{\circ}$, $\alpha+180^{\circ}$}, covering 360 degrees of the human head. Elevation is 0 for all images. We opt for six views as the optimal balance - fewer views compromise detail quality while more views become computationally prohibitive for full head reconstruction. An ablation study comparing four, six, and eight views is presented in Sec. 5.2.
Multi-view Attention. To ensure the consistency of the generated novel views, we use a multi-view attention mechanism to facilitate information propagation and implicitly encode multi-view dependencies. Our attention module encourages multi-view consistency by extending the 2D self-attention mechanism to 3D and enabling interactions across views. Instead of treating each view independently, we apply self-attention across all views simultaneously, allowing information to be shared between them. Specifically, we start with an input tensor of shape $B\!\times\!V\!\times\!H\!\times\!W\!\times\!C$, where $B$ is the batch size, $V$ is the number of views, $H$ and $W$ denote the spatial resolution of the intermediate feature maps, and $C$ is the number of feature channels. We reshape this tensor to $B\!\times\!VHW\!\times\!C$, treating all spatial locations across all views as a unified sequence of tokens for self-attention. This design allows the model to learn multi-view correlations by sharing information across views within the attention layers, enabling it to generate consistent RGB images. We provide an ablation study on the multi-view attention mechanism in the supplementary material.
Input View Reconstruction. During Training, we enforce the first generated view to share the same camera with the input image. In other words, we reconstruct the input view in the view generation process. We find this approach, combined with the multi-view attention mechanism, significantly outperforms the alternative strategy of generating only novel views, which tends to overfit to synthetic training identities and compromises generalization capability as we will show in Sec. 5.1 and the supplementary material.
3.3 Multi-view to 3D Gaussians Reconstruction
Transformer-based Reconstructor. We choose pixel-aligned 3D Gaussians as the final 3D representation. Each Gaussian $G_{i}$ is defined by position $p_{i}$, scale $s_{i}$, orientation $q_{i}$, opacity $\alpha_{i}$, and color features $c_{i}$. We use a transformer-based reconstructor $f_{G}$ to obtain 3D Gaussians from generated multi-view images $x_{0}^{1:N}$ and their corresponding Plücker ray coordinates [44] $\textsc{P}^{1:N}$:
$\{G_{i}\}_{i=1}^{NHW}\!=\!\{p_{i},s_{i},q_{i},\alpha_{i},c_{i}\}_{i=1}^{NHW}\!=\!f_{G}(x_{0}^{1:N},\textsc{P}^{1:N}),$ (3)
Our $f_{G}$ is a large reconstruction model [24, 74] which follows the implementation of GS-LRM [74]: the $N$ multi-view images are concatenated with their Plücker ray coordinates computed from the camera intrinsic and extrinsic parameters for pose conditioning. Then, the inputs are patchified by dividing the per-view feature map into non-overlapping patches with a patch size of $p$. Each 2D patch is then flattened into a 1D vector. Finally, a linear layer L is utilized to map the 1D vectors to image patch tokens:
$\{\textsc{T}_{j}^{n}\}_{j=1,2,\dots,\frac{HW}{p^{2}}}=\textit{L}(\texttt{Patchify}_{p}(\texttt{Concat}(\textsc{I}^{n},\textsc{P}^{n}))).$ (4)
Where $\{\textsc{T}_{j}^{n}\}$ denotes the set of patch tokens for image $n$, totaling $\frac{HW}{p^{2}}$ tokens per image. The set of multi-view image tokens $\{\textsc{T}_{j}^{n}\}$ are concatenated and processed through a chain of transformer blocks. Each transformer block is equipped with residual connections [20] and consists of Pre-LayerNorm [2], multi-head Self-Attention [59] and MLP. Later, the output tokens from the transformer are decoded into Gaussian parameters using a single linear layer. Then, the Gaussian parameters are unpatchified into $p^{2}$ Gaussians. Finally, we end up with $HW$ Gaussians for each view, where pixel encodes one 3D Gaussian.
Two-stage Training. We find that training the transformer-based reconstructor solely on synthetic human head data leads to inferior texture details when applied to real-world images (see ablation study in Fig. 12). We suspect this limitation arises because the synthetic datasets lack geometric diversity. To address this, we propose a two-stage training approach in which the reconstructor is pre-trained on diverse object data [12] and subsequently fine-tuned using synthetic head data. The pre-training stage enables the reconstructor to learn a broad prior of diverse geometric structures, yielding more detailed and clearer textures in delicate facial regions such as the eyes, nose, and ears. The fine-tuning process then imparts specific knowledge of head structure, producing smoother and more realistic results. During training, we randomly select four input views to reconstruct a total of eight views, four input and four novel views. Following [74], we optimize the model using a combination of MSE and perceptual losses. During inference, the reconstructor processes the six-view outputs from multi-view diffusion model to reconstruct the head.
3.4 Real-world Image Inference
For inference on real-world images, since their intrinsic parameters are unknown, we adopt a camera fov of $50^{\circ}$, same as during training. To ensure plausible outputs, we first apply an MTCNN face detector to estimate the face’s size and center. The image is then resized and cropped/extended to match the average face size and center computed from the training data. We find this alignment compensates for the unknown intrinsic parameters well, ensuring plausible reconstruction results.